OpenStudy (anonymous):

There are 6 girls and 7 boys in a class. A team of 10 players is to be selected from the class. How many different combinations of players are possible? 10 13 130 286

OpenStudy (anonymous):

@amoodarya

OpenStudy (triciaal):

total of 6 plus 7; need to choose 10

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

i CAN'T get this wrong

OpenStudy (anonymous):

i wont pass if i do

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

@superman36

Vocaloid (vocaloid):

combination formula number of choices = nCr = n!/(r!*(n-r)!) where n = total number of people (13) and r = the number of people on the team = 10

OpenStudy (anonymous):

thank u so much!

OpenStudy (triciaal):

welcome

OpenStudy (anonymous):

thanks to u too :) @triciaal

OpenStudy (anonymous):

FUTURE PLATO PEOPLE! its not 10!!!

OpenStudy (triciaal):

does this mean you did not answer 286?

OpenStudy (anonymous):

yea...

OpenStudy (anonymous):

i put 10

OpenStudy (anonymous):

both of u told me it was 10!

OpenStudy (triciaal):

why on earth did you do that after saying how important it was to get it right!

OpenStudy (anonymous):

u guys told me it was 10! lol

OpenStudy (triciaal):

NO I DID NOT!

OpenStudy (anonymous):

yea u did! xD

OpenStudy (triciaal):

you have 13 to choose 10

OpenStudy (anonymous):

so what was the answer!?

OpenStudy (triciaal):

you are not reading slow enough!

OpenStudy (anonymous):

286?

OpenStudy (triciaal):

yes as shown by the other response put in on the calculator using nCr 13C10

OpenStudy (anonymous):

idk how to work that! omg

OpenStudy (triciaal):

so sorry you got it wrong but now you know to slow down,

OpenStudy (anonymous):

ik lol im gonna read super slow next time

OpenStudy (anonymous):

stupid test xD

OpenStudy (amoodarya):

\[\left(\begin{matrix}6+7 \\ 10\end{matrix}\right)=\left(\begin{matrix}13 \\ 10\end{matrix}\right)=\\\left(\begin{matrix}13 \\ 3\end{matrix}\right)=\frac{13*12*11}{3.2.1}=13*2*11\]